### v2

@@ 1 Introduction
Nearly all problems in astrodynamics involve the force of gravity. Be it in trajectory optimization, spacecraft rendezvous, orbit determination, or other problems, gravity often plays a significant — if not dominant — role in the system dynamics. The mere ubiquity of this force is a testament to its significance, yet despite this, there exist surprisingly few ways to represent this force to high accuracy. The construction of high-fidelity gravity models is henceforth referred to as the gravity modeling problem, and currently there exists no universally adopted solution. Some gravity models perform well for large planetary bodies, but break down when modeling objects with exotic geometries like asteroids and comets. Other models can handle irregular shapes but require assumptions and come with high computational cost. The one commonality is that all existing gravity models come with their own unique pitfalls that prevent standardization across the community. Consequently, researchers are continuing to design new solutions to the gravity modeling problem with hopes of one day finding a universal model. One encouraging vein of research explores the use of neural networks and scientific machine learning to learn gravity models free of these past difficulties.
@@ 2.1 Analytic Models
While these models are effective at representing planetary oblateness, they struggle to model the remaining gravitational perturbations like mountain ranges, tectonic plate boundaries, and craters. These discontinuous features are notoriously hard to represent using periodic basis functions — requiring the superposition of hundreds of thousands of high-frequency harmonics to overcome the 3D equivalent of Gibbs phenomenon hewittGibbsWilbrahamPhenomenonEpisode1979 . These higher frequencies introduce an $\mathcal{O}(n^{2})$ computational and memory cost martinGPGPUImplementationPines2020 , and the regression of these harmonics is especially difficult due to their rapidly fading observability. High-order spherical harmonic models also can diverge when evaluated within the sphere that bounds all mass elements (Brillouin sphere) due to the $(R/r)^{l}$ term in the series expansion. While these effects are negligible for near-spherical planets or moons, they can become problematic in small-body settings where objects can exhibit highly non-spherical geometries wernerExteriorGravitationPolyhedron1997 .
@@ 2.2 Machine Learning Models
Finally, in 2022 Reference 25 introduced GeodesyNets as a candidate solution to the gravity modeling problem. GeodesyNets are neural density fields — close relatives to the popular neural radiance fields mildenhallNeRFRepresentingScenes2022 — which uses neural networks to learn a density function for every point in a 3D volume. Once learned, these density fields can be integrated to yield a gravitational acceleration or potential. This work is an exciting new class of machine learning model capable of achieving high accuracy; albeit the cost of the numerical integration of densities make it cumbersome for simulation or on-board purposes. These models also have more computationally expensive training processes than other neural network approaches.
@@ 3 PINN-GM-III
This data can be collected in one of two ways: 1) It can be generated synthetically from pre-existing high-fidelity gravity models like EGM-2008 for Earth pavlisDevelopmentEvaluationEarth2012 or a high-fidelity polyhedral model for asteroids gaskellGaskellErosShape2021 , or 2) the data can be estimated on-board without any prior knowledge of the field. The latter can be accomplished using estimation techniques like dynamic model compensation leonardGravityErrorCompensation2013 to simultaneously estimate spacecraft position and accelerations in-situ as is shown in Ref. martinPINNFilter2022 .
@@ 3.3 Improve Numerics by Learning a Proxy to the Potential
Another common problem of past machine learning gravity models is numerical clipping. These models use the same matrix operations (i.e. the neural network) to predict both very large and very small values of the potential depending on the test point’s altitude. Unfortunately, these scales can lead to difficult numerics. When the altitude exceeds some critical threshold, the predictions grow too small, and the network prematurely clips their value to zero. This is problematic for both model training and inference — prematurely capping the maximum altitude for which these models are usable.
@@ 3.5 Leveraging Preexisting Gravity Information into PINN-GM Solution
where $\hat{U}_{\text{LF}}$ is the weighted low fidelity analytic model with $w_{\text{LF}}=H(r,R^{*},k^{*})$, $R^{*}=1+e$, and $k=0.5$, where $e$ is the eccentricity of the body computed via $e=\sqrt{1-b^{2}/a^{2}}$. This ensures that at high altitude, where the analytic approximation is most accurate, it fully contributes to the final learned solution, but at lower altitudes, the analytic contribution has a reduced weight to the final solution. A visualization of the various weighting functions used in the PINN-GM-III is shown in Figure 3, highlighting where the different parts of the model are fully activated and deactivated.
@@ 3.6 Visualizing Modifications’ Effects on Model Performance
To visualize how the proposed modifications affect model accuracy, six PINN-GM are trained. The first model corresponds with the original PINN-GM-II, and the subsequent models sequentially add one of the modification discussed above. The PINN-GM are trained with a 5,000 point dataset uniformly distributed between the surface of the asteroid Eros to an altitude of 15R. The corresponding acceleration error of each model is reported within the training distribution in Figure 4(a) and outside the training distribution in Figure 4(b).
Figure 4(a) illustrates that these design modifications consistently improve the modeling accuracy of the PINN within the bounds of the training distribution. The PINN-GM-II had errors approaching 1% at the high 15R altitudes; however this decreases to 0.2%, 0.05%, 0.01%, 0.03%, and 0.001% for each of the respective modifications. Notably the transition to the new loss function (II) does decrease performance near the surface, but this is to be expected given the redistribution of modeling priorities that now balance accuracy at both low- and high-altitudes.
Figure 4(b) illustrates the stabilizing effect of these modifications on extrapolation error — quantifying prediction error outside of the training bounds out to 100R. As is shown in blue, the PINN-GM-II entirely diverges after leaving the bounds of the training data at 15R; however, the added design modifications stabilize performance and prevent divergence. As was expected, the error in the high-altitude limit hits a numerical plateau when only the features (I) and percent loss (II) modifications are included; however, when the proxy potential (III) is used, that plateau lowers by over an order-of-magnitude. The boundary condition modification (IV) eliminates this plateau — albeit inducing a slight penalty near the transition point due to changes in the gradient — however, the addition of the low-fidelity point mass potential (V) eliminates this penalty and further reduces the error within the bounds of the training data. It should be noted that there exists a fundamental numerical limit to the percent error at approximately 0.001% where the differences between model prediction and the ground truth exists at the limit of machine precision.
@@ 4 Benchmarking Suite
While the proposed modifications yield improvements as presented in Section 3.6, it remains important to characterize these gravity models from many different perspectives. The lack of comprehensive and standard performance benchmarks for these models are part of the reason these failure cases had not yet been identified. To eliminate this possibility moving forward and to further scrutinize the proposed PINN-GM-III, six new evaluation metrics are proposed. These metrics characterize gravity model accuracy within and beyond the training distribution, aiming to provide both coarse and fine measures of performance in different orbital regimes. These metrics are explained in detail in Section 4.1 and are used to characterize the performance of a trained PINN-GM-III on a heterogeneous density asteroid in Section 4.2.
@@ 4.1 Metrics
The second, third, and fourth metrics investigate the generalization of the model across a range of altitudes both within and beyond the training bounds. Explicitly, the mean acceleration error is evaluated as a function of altitude, and divided into three testing regimes: interior, exterior, and extrapolation. The interior metric assesses error within the bounding sphere of radius R. The exterior metric investigates the error out to the maximum altitude of the training dataset. Finally, the extrapolation metric measures the error exclusively outside the training dataset — specifically reaching altitudes 10 times larger than the maximum altitude represented in the training set. For every unit of radius, 500 samples are distributed uniformly in altitude to produce the test set.
@@ 4.2 Experiment
These metrics are used to evaluate of a PINN-GM-III trained on data generated from a synthetic, heterogeneous density asteroid modeled after 433-Eros. Heterogeneous asteroids provide an especially challenging scenario for gravity models, as their internal density distributions are not directly observable. Some asteroids contain over- and under-dense regions within their interior, or may have been formed by two asteroids merging together. While some heuristic methods have been proposed to estimate these asteroids’ internal densities takahashiMorphologyDrivenDensity2014 which can be used in heterogeneous forms of the polyhedral model takahashiMorphologyDrivenDensity2014 ; the more common practice is to simply proceed with a constant density assumption.
To induce this heterogeneous density body, two small mass inhomogeneities are placed inside the asteroid 433-Eros. In one hemisphere, a mass element is added, and in the other hemisphere, a mass element is removed. Each mass element contains 10% of the total mass of the asteroid, and they are symmetrically displaced along the x-axis by 0.5R (see Figure 5(a)). The gravitational contributions of these mass elements are superimposed onto the gravity field of a constant density polyhedral model to form the simulated ground truth. This choice emulates the gravity field of a single body formed by two merged asteroids of different characteristic densities. The choice to make each mass element ±$\pm 10$% of the total mass is motivated based on literature with similar candidate density distributions kanamaruEstimationInteriorDensity2019 .
Using this heterogeneous density model, 90,000 position and acceleration data are sampled uniformly around the body from 0-10R. An additional 200,700 points are sampled on the surface of the asteroid — corresponding to a data point on every facet of the asteroid shape model. Together these approximately $300,000$ data points constitute a “best case” training set for the PINN-GM-III. These conditions are purposefully used to evaluate the upper-bound performance of the model. The next section will explore model performance under more realistic and stressful data conditions.
Using this best case data, a PINN-GM-II and a PINN-GM-III are trained. Both networks have six hidden layers with 32 nodes per layer which corresponds to approximately 5,300 learnable parameters. The default hyperparameters used to train these models can be found in Appendix A alongside the ablation studies used to select them in Appendix B. Once trained, these PINN-GMs are evaluated using the aforementioned metrics and compared to a constant-density polyhedral model as shown in Figure 6.
@@
The generalization metrics are shown at the top of Figure 6, highlighting model performance across the three altitude regimes. The constant density polyhedral model produces the highest error both on the interior ($<$R) and exterior (1-10R) metrics with errors averaging around 10% within Brillouin sphere and above 1% error outside of it. In the extrapolation regime, the polyhedral model error slowly decreases as the model tends towards a point mass approximation. In comparison, the PINN-GM-II produced average errors of 0.5% inside the bounding sphere, and 1% outside the bounding sphere, with errors slowly increasing at higher altitudes as a result of its original absolute error loss function. In the extrapolation regime, the PINN-GM-II entirely diverges for the reasons discussed in Section 3. In contrast to these models, the PINN-GM-III maintains lower errors across the interior and exterior distributions, averaging less than 0.5% error within the bounding sphere and approximately 0.005% error outside of it. As the model transitions into the extrapolation regime, the Heaviside function kicks in to enforce the point mass boundary condition — helping the network converge to the same limit as the polyhedral model.
The planes and surface metrics are shown in the middle rows of Figure 6. The polyhedral gravity model produced 6.5% average acceleration error on the planes, the PINN-GM-II reduces this to 0.4%, and the PINN-GM-III to 0.07%. The errors are largest near the surface for models, averaging at 23% for the constant density polyhedral model, 0.2% for the PINN-GM-II, and 0.18% for the PINN-GM-III.
@@ 5 Comparative Study
In this section, a comparative study is performed to evaluate the PINN-GM-III against other popular gravity models. Explicitly, a point mass (PM), spherical harmonic (SH), mascon, polyhedral, extreme learning machine (ELM), traditional neural network (TNN), PINN-GM-I, and PINN-GM-II model are each regressed and then evaluated using the aforementioned metrics. Each model is fit eight times, permuting the data quantity, data uncertainty, and the size of each model. [^1] Note: The point mass model only has four permutations, as it does not have a “large” option and the polyhedral model only has two permutations, large and small. This is because polyhedral models are regressed in an entirely unrelated fashion to the other models — relying primarily on image data — and their sensitivity to data quantity and quality can not be adequately compared through this proposed experiment. Therefore, only the effect of model size is considered, and the performance can be considered an upper-bound. These conditions purposefully test model robustness under more stressful data conditions and different model sizes.
@@
PINN III S 50000 0 29 0.36 2.6 0.060 0.13 18 0.056
PINN III L 50000 0 31 1.2E-2 0.11 1.2E-3 1.1 3.3 8.4E-3
MASCONS L 50000 0 41 1.2E-2 0.13 5.4E-4 3.4E-4 33 2.7
PINN III S 500 0 64 1.5 8.6 0.35 0.27 31 1.1
PINN II S 50000 0 70 0.58 1.7 0.63 1.5E5 13 0.27
PINN II L 50000 0 73 1.4E-2 0.059 1.4E-2 1.4E15 2.8 9.6E-3
PINN III L 50000 10 76 3.4 3.7 2.0 0.28 13 6.3
MASCONS L 500 0 84 0.74 11 2.1E-4 9.8E-8 310 2.7
PINN III S 50000 10 92 3.9 8.3 2.0 0.34 35 3.0
MASCONS L 50000 10 94 1.3 9.3 0.16 0.15 630 2.4
POLY S 94 6.2 16 3.3 0.34 23 1.9
PINN II S 50000 10 95 2.1 3.2 2.9 1.8E5 15 1.4
PINN III L 500 0 98 1.6 17 0.21 0.37 51 0.41
PINN II L 50000 10 101 1.4 2.7 2.2 4.1E14 8.9 0.80
PINN II L 500 0 102 0.70 5.5 0.22 5.0E12 26 0.20
POLY L 110 6.5 16 3.5 0.35 23 2.3
MASCONS S 50000 0 119 3.1 4.4 2.8 2.7 41 8.2
PINN II S 500 0 122 2.5 8.2 2.9 3.2E5 30 3.0
MASCONS S 50000 10 128 3.3 4.6 3.0 3.0 42 9.0
MASCONS S 500 0 137 2.7 12 1.9 1.8 150 6.4
PINN I L 50000 0 140 10 4.2 8.3 1.4E5 8.8 11
PINN III L 500 10 143 6.1 27 3.0 0.36 49 6.4
PINN II S 500 10 150 5.8 12 6.6 4.3E5 32 2.5
PINN III S 500 10 157 11 58 3.4 0.31 82 22
PINN I L 50000 10 161 14 6.1 30 1.8E5 17 13
PINN I S 50000 0 165 14 8.3 38 1.3E5 31 25
PINN I S 50000 10 169 16 10 38 1.4E5 33 4.8
PINN II L 500 10 180 5.9 13 5.0 4.3E11 32 7.1
PINN I L 500 10 183 31 17 49 45000 33 32
SH S 500 10 189 25 160 5.0 0.35 1500 4.4
PINN I S 500 10 195 33 19 70 45000 37 36
SH S 500 0 198 26 160 5.0 0.35 1500 4.4
PINN I S 500 0 207 38 24 75 95000 42 47
PINN I L 500 0 210 49 38 100 1.1E5 38 55
SH S 50000 10 213 240 2000 5.0 0.35 32000 4.4
SH S 50000 0 214 230 2000 5.0 0.35 31000 4.4
PM - 500 0 215 50 43 50 50 70 120
MASCONS S 500 10 216 34 98 29 28 480 56
PM - 500 10 219 50 43 51 51 70 120
PM - 50000 10 223 53 45 53 53 70 120
PM - 50000 0 227 53 45 53 53 70 120
SH L 50000 0 240 inf 8.4E99 5.0 0.35 3.4E117 4.4
SH L 50000 10 240 inf 8.9E99 5.0 0.35 3.5E117 4.4
SH L 500 10 245 inf 2.5E97 5.0 0.35 1.1E115 4.4
SH L 500 0 251 inf 9.8E96 5.0 0.35 2.1E115 4.4
ELM L 500 0 258 76 97 300 8.0E7 99 65
ELM S 500 0 258 74 98 250 2.0E7 99 66
ELM L 500 10 260 78 97 310 8.4E7 99 65
ELM S 500 10 260 76 98 250 2.1E7 99 66
TNN S 500 10 273 770 350 1200 66000 240 640
TNN L 500 10 278 510 1100 800 56000 880 500
TNN S 50000 0 295 610 1100 1800 1.7E5 1100 580
TNN S 500 0 301 1300 1100 2600 1.5E5 990 450
TNN L 50000 10 308 630 1100 2300 2.2E5 1200 800
TNN L 50000 0 310 550 1100 1800 4.5E5 1700 640
TNN S 50000 10 310 640 1100 2400 2.2E5 1100 760
TNN L 500 0 312 1300 1300 4800 1.6E5 1100 780
ELM L 50000 10 314 9100 990 53000 2.6E9 510 47000
ELM S 50000 10 318 9200 1000 54000 1.8E8 520 1.9E6
MASCONS L 500 10 319 4.2E11 7.7E12 1.6E8 0.58 1.7E14 inf
ELM L 50000 0 324 9500 1000 56000 2.6E9 540 51000
ELM S 50000 0 328 9700 1100 57000 1.9E8 550 2.0E6

### v3

@@ 1 Introduction
Nearly all problems in astrodynamics involve the force of gravity. Be it in trajectory optimization, spacecraft rendezvous, orbit determination, or other problems, gravity often plays a significant — if not dominant — role in the system dynamics. The ubiquity of this force is a testament to its significance, yet despite this, there exist surprisingly few ways to represent this force to high accuracy. The construction of high-fidelity gravity models is henceforth referred to as the gravity modeling problem, and currently there exists no universally adopted solution. Some gravity models perform well for large planetary bodies, but break down when modeling objects with exotic geometries like asteroids and comets. Other models can handle irregular shapes but require assumptions and come with high computational cost. The one commonality is that all existing gravity models come with their own unique pitfalls that prevent standardization across the community. Consequently, researchers are continuing to design new solutions to the gravity modeling problem with hopes of one day finding a universal model. One encouraging vein of research explores the use of neural networks and scientific machine learning to learn gravity models free of these past difficulties.
@@ 2.1 Analytic Models
While these models are effective at representing planetary oblateness, they struggle to model the remaining gravitational perturbations like mountain ranges, tectonic plate boundaries, and craters. These discontinuous features are notoriously hard to represent using periodic basis functions — requiring the superposition of hundreds of thousands of high-frequency harmonics to overcome the 3D equivalent of Gibbs phenomenon hewittGibbsWilbrahamPhenomenonEpisode1979 . These higher frequencies introduce an $\mathcal{O}(n^{2})$ computational and memory cost martinGPGPUImplementationPines2020 , and the regression of these harmonics is especially difficult due to their rapidly fading observability. High-order spherical harmonic models also can diverge when evaluated within the sphere that bounds all mass elements (Brillouin sphere). While these effects are negligible for near-spherical planets or moons, they can become problematic in small-body settings where objects can exhibit highly non-spherical geometries wernerExteriorGravitationPolyhedron1997 .
@@ 2.2 Machine Learning Models
Finally, in 2022 Reference 25 introduced GeodesyNets as a candidate solution to the gravity modeling problem. GeodesyNets are neural density fields — close relatives to the popular neural radiance fields mildenhallNeRFRepresentingScenes2022 — which use neural networks to learn a density function for every point in a 3D volume. Once learned, these density fields can be numerically integrated to yield a gravitational acceleration or a value of the potential. This work is an exciting new class of machine learning model capable of achieving high accuracy given sufficient quantities of data. These models also behave more reliably at high-altitudes as the networks are not required to predict densities beyond a unit volume; albeit the cost required to perform the numerical integration can be relatively high — requiring evaluation of 30,000 to 300,000 quadrature points per prediction izzoGeodesyIrregularSmall2022 — which can make these models cumbersome to evaluate and train. Finally, a key feature of these models is their ability to estimate an internal density profile of an asteroid — a capability presumed available to PINNs by evaluating Poisson’s equation inside the body — but such capabilities are not the focus of this work.
@@ 3 PINN-GM-III
This data can be collected in one of two ways: 1) It can be generated synthetically from pre-existing high-fidelity gravity models like EGM-2008 for Earth pavlisDevelopmentEvaluationEarth2012 or a high-fidelity polyhedral model for asteroids gaskellGaskellErosShape2021 , or 2) the data can be estimated on-board without any prior knowledge of the field. The latter can be accomplished using estimation techniques like dynamic model compensation leonardGravityErrorCompensation2013 to simultaneously estimate spacecraft position and accelerations in-situ as is shown in Ref. martinPINNFilter2022 . For this study, all data will be synthetically generated and evaluation on real data is left for future work.
@@ 3.3 Improve Numerics by Learning a Proxy to the Potential
Another common problem of past machine learning gravity models is numerical clipping. These models use the same matrix operations (i.e. the neural network) to predict both very large and very small values of the potential depending on the test point’s altitude, which can pose difficult numerics. When the altitude exceeds some critical threshold, the predictions grow too small, and the network prematurely clips their value to zero. This is problematic for both model training and inference — prematurely capping the maximum altitude for which these models are usable.
@@ 3.5 Leveraging Preexisting Gravity Information into PINN-GM Solution
where $\hat{U}_{\text{LF}}$ is the weighted low fidelity analytic model with $w_{\text{LF}}=H(r,R^{*},k^{*})$, $R^{*}=0$, and $k=0.5$. This ensures that at high altitude, where the analytic approximation is most accurate, it fully contributes to the final learned solution, but at lower altitudes, the analytic contribution has a reduced weight to the final solution. A visualization of the various weighting functions used in the PINN-GM-III is shown in Figure 3, highlighting where the different parts of the model are fully activated and deactivated.
@@ 3.6 Visualizing Modifications’ Effects on Model Performance
To visualize how the proposed modifications affect model accuracy, six PINN-GM are trained. The first model corresponds with the original PINN-GM-II, and the subsequent models sequentially add the proposed modifications. The PINN-GM are trained with a 5,000 point dataset uniformly distributed between the surface of the asteroid Eros to an altitude of 15R. The corresponding acceleration error of each model is reported within the training distribution in Figure 4(a) and outside the training distribution in Figure 4(b).
Figure 4(a) illustrates that these design modifications consistently improve the modeling accuracy of the PINN within the bounds of the training distribution. The PINN-GM-II had errors approaching 1% at the high 15R altitudes; however this decreases to 0.2%, 0.05%, 0.01%, 0.03%, and 0.001% for each of the respective modifications. Notably the transition to the new loss function (II) does reduce accuracy near the surface, but this is to be expected given the redistribution of modeling priorities that now balance accuracy at both low- and high-altitudes.
Figure 4(b) illustrates the stabilizing effect of these modifications on extrapolation error — quantifying prediction error outside of the training bounds out to 100R. As is shown in blue, the PINN-GM-II entirely diverges after leaving the bounds of the training data at 15R; however, the design modifications stabilize performance and prevent this divergence. When only the features (I) and percent loss (II) modifications are included, the error in the high-altitude limit hits a numerical plateau as expected; however, the inclusion of the proxy potential (III) lowers that plateau by over an order-of-magnitude. The boundary condition modification (IV) eliminates this plateau — albeit inducing a small penalty near the transition point due to changes in the gradient. Finally, the addition of the low-fidelity point mass potential (V) eliminates this penalty and further reduces the error at higher within the bounds of the training data.
@@ 4 Benchmarking Suite
The lack of comprehensive and standard performance benchmarks for gravity models are part of the reason these failure cases had not yet been identified. To eliminate this possibility moving forward and to further scrutinize the proposed PINN-GM-III, six new evaluation metrics are proposed. These metrics characterize gravity model accuracy within and beyond the training distribution, aiming to provide both coarse and fine measures of performance in different orbital regimes. These metrics are explained in detail in Section 4.1 and are used to characterize the performance of a trained PINN-GM-III on a heterogeneous density asteroid in Section 4.2.
@@ 4.1 Metrics
The second, third, and fourth metrics investigate the generalization of the model across a range of altitudes both within and beyond the training bounds. Explicitly, the mean acceleration error is evaluated as a function of altitude and divided into three testing regimes: interior, exterior, and extrapolation. The interior metric assesses error within the bounding sphere of radius R. The exterior metric investigates the error out to the maximum altitude of the training dataset. Finally, the extrapolation metric measures the error exclusively outside the training dataset — specifically reaching altitudes 10 times larger than the maximum altitude represented in the training set. For every unit of radius, 500 samples are distributed uniformly in altitude to produce the test set.
@@ 4.2 Experiment
These metrics are used to evaluate a PINN-GM-III trained on data generated from a synthetic, heterogeneous density asteroid modeled after 433-Eros. Heterogeneous asteroids provide an especially challenging scenario for gravity models, as their internal density distributions are not directly observable. Some asteroids contain over- and under-dense regions within their interior, or may have been formed by two asteroids merging together. While some heuristic methods have been proposed to estimate these asteroids’ internal densities — which can then be used in heterogeneous forms of the polyhedral model takahashiMorphologyDrivenDensity2014 — the more common practice is to simply proceed with a constant density assumption.
To induce this heterogeneous density body, two small mass inhomogeneities are placed inside the asteroid. In one hemisphere, a mass element is added, and in the other hemisphere, a mass element is removed. Each mass element contains 10% of the total mass of the asteroid, and they are symmetrically displaced along the x-axis by 0.5R (see Figure 5(a)). The gravitational contributions of these mass elements are superimposed onto the gravity field of a constant density polyhedral model to form the simulated ground truth. This choice emulates the gravity field of a single body formed by two merged asteroids of different characteristic densities. The choice to make each mass element ±$\pm 10$% of the total mass is motivated based on literature with similar candidate density distributions kanamaruEstimationInteriorDensity2019 .
Using this heterogeneous density model, 90,000 position and acceleration data are sampled uniformly around the body from 0-10R. An additional 200,700 points are sampled on the surface of the asteroid — corresponding to a data point on every facet of the asteroid shape model. Together these approximately $300,000$ data points constitute a “best case” training set for the PINN-GM-III used to evaluate the upper-bound performance of the model. Subsequent sections explore model performance under more realistic and stressful data conditions.
Using this data, a PINN-GM-II and a PINN-GM-III are trained. Both networks have six hidden layers with 32 nodes per layer which corresponds to approximately 5,300 learnable parameters. The default hyperparameters used to train these models can be found in Appendix A alongside the studies used to select them in Appendix B. Once trained, these PINN-GMs are evaluated using the aforementioned metrics and compared to a constant-density polyhedral model as shown in Figure 6.
@@
The generalization metrics are shown at the top of Figure 6, highlighting model performance across the three altitude regimes. The constant density polyhedral model produces the highest error on the interior and exterior metrics, averaging 10% and 1% respectively. In the extrapolation metric, the polyhedral error decreases as the model begins to behave like a point mass approximation. In comparison, the PINN-GM-II produces lower average errors on the interior and exterior metrics — 0.5% and 1% — however the candidate pitfalls described in Section 3 also become apparent. The former RMS loss function yields a monotonically increasing error as a function of altitude within the bounds of the training data, and the model diverges outside of it. In contrast, the PINN-GM-III maintains the lowest errors across all three regimes, averaging less than 0.5% and 0.005% error on the interior and exterior metrics, and maintains stability in the extrapolation regime due to the hyperbolic tangent function which enforces the point mass boundary conditions.
The planes and surface metrics are shown in the middle rows of Figure 6. The polyhedral gravity model produced an average acceleration error of 6.5% on the planes, the PINN-GM-II reduces this to 0.4%, and the PINN-GM-III to 0.07%. The errors are largest near the surface for all models, averaging at 23% for the constant density polyhedral model, 0.2% for the PINN-GM-II, and 0.18% for the PINN-GM-III.
@@ 5 Comparative Study
A comparative study is performed to evaluate the PINN-GM-III against other popular gravity models. Explicitly, a point mass (PM), spherical harmonic (SH), mascon, polyhedral, extreme learning machine (ELM), traditional neural network (TNN), PINN-GM-I, PINN-GM-II, and geodesyNet model are each regressed and then evaluated using the aforementioned metrics. Each model is fit eight times, permuting the data quantity, data uncertainty, and the size of each model. [^1] Note: The point mass model only has four permutations, as it does not have a “large” option and the polyhedral model only has two permutations, large and small. This is because polyhedral models are regressed in an entirely unrelated fashion to the other models — relying primarily on image data — and their sensitivity to data quantity and quality can not be adequately compared through this proposed experiment. Therefore, only the effect of model size is considered, and the performance can be considered an upper-bound. These conditions purposefully test model robustness under more stressful data conditions and different model sizes.
@@
MASCONS L 50000 0 14 1.2E-2 0.13 5.4E-4 3.4E-4 8.3 1.7E-3
PINN III S 50000 0 29 0.17 1.2 2.8E-2 0.065 11 4.2E-2
PINN III L 50000 0 34 1.3E-2 0.099 1.7E-3 1.3 3.2 2.7E-3
PINN II L 50000 0 84 4.7E-2 0.080 0.070 3.2E14 2.5 2.8E-2
PINN III S 500 0 88 2.2 18 0.51 0.13 35 0.89
PINN III L 50000 10 89 3.4 3.7 2.0 0.28 13 6.3
PINN II S 50000 0 97 0.22 0.57 0.31 2.1E7 6.2 0.11
PINN III L 500 0 98 1.6 17 0.23 0.41 52 0.67
MASCONS L 500 0 101 0.82 14 2.5E-4 9.8E-8 350 1.3E-7
POLY S 110 6.2 16 3.3 0.34 23 1.9
MASCONS L 50000 10 113 1.3 9.3 0.16 0.15 610 0.37
PINN II L 500 0 115 0.69 5.1 0.22 3.2E12 25 0.18
MASCONS S 50000 0 117 3.1 4.4 2.8 2.7 26 12
POLY L 119 6.5 16 3.5 0.35 23 2.3
PINN II L 50000 10 122 1.4 2.7 2.2 4.4E14 8.8 0.68
MASCONS S 50000 10 124 3.3 4.6 3.0 3.0 26 13
MASCONS S 500 0 135 2.7 12 1.9 1.8 71 6.7
PINN III L 500 10 140 6.2 23 3.1 0.36 45 6.0
PINN II S 50000 10 140 1.5 2.9 1.8 1.1E8 9.4 6.4
PINN II S 500 0 152 1.8 7.8 1.7 5.4E6 32 3.2
PINN III S 500 10 159 8.2 51 3.3 0.36 76 5.0
PINN III S 50000 10 160 8.1 51 2.3 0.32 79 17
PINN I L 50000 0 174 10 4.3 8.4 1.6E5 8.4 12
PINN II L 500 10 179 5.7 13 4.8 5.6E11 32 5.4
PINN II S 500 10 182 6.4 13 6.5 8.6E6 33 2.7
PINN I L 50000 10 183 13 5.4 16 1.8E5 14 7.0
SH S 50000 0 202 24 370 0.45 0.22 1900 1.1
PINN I S 50000 10 202 14 9.5 31 1.6E5 31 20
PINN I S 50000 0 203 16 8.6 38 1.2E5 33 17
SH S 50000 10 211 24 320 0.53 0.29 2200 1.4
PM - 500 0 211 50 43 50 50 70 110
PM - 500 10 215 50 43 51 51 70 110
SH S 500 0 217 17 340 1.2 0.86 950 0.97
PM - 50000 10 220 53 45 53 53 70 110
PINN I L 500 10 221 29 16 48 46000 34 22
PM - 50000 0 224 53 45 53 53 70 110
SH S 500 10 227 27 580 1.6 0.74 1500 2.4
PINN I S 500 10 235 33 21 69 46000 37 31
SH L 50000 0 239 inf 1.3E99 0.45 0.22 8.8E116 1.1
PINN I S 500 0 242 41 24 75 93000 41 42
GEONET L 500 0 245 83 84 83 83 87 160
SH L 50000 10 247 inf 1.8E99 0.53 0.29 1.2E117 1.4
MASCONS S 500 10 251 34 98 29 28 460 64
GEONET L 50000 0 252 87 87 87 87 87 160
GEONET L 500 10 255 87 90 86 86 93 160
GEONET L 50000 10 260 87 87 87 87 87 160
GEONET S 500 0 273 99 99 99 99 99 180
SH L 500 10 275 inf 2.5E97 5.0 0.35 1.0E115 4.4
SH L 500 0 278 inf 9.8E96 5.0 0.35 2.1E115 4.4
PINN I L 500 0 280 52 41 100 1.1E5 47 89
GEONET S 50000 10 281 99 99 99 99 99 180
GEONET S 500 10 288 99 99 99 99 99 180
GEONET S 50000 0 292 99 99 99 99 99 180
ELM L 500 0 308 76 97 300 8.0E7 99 65
ELM L 500 10 308 78 97 310 8.4E7 99 65
ELM S 500 0 315 74 98 250 2.0E7 99 66
ELM S 500 10 315 76 98 250 2.1E7 99 66
MASCONS L 500 10 369 3.6E11 5.5E12 1.7E8 0.63 1.4E14 inf
TNN L 500 10 408 570 1200 810 56000 1100 420
TNN S 500 0 409 1300 1100 2600 1.5E5 990 450
TNN S 50000 0 410 630 1100 1900 1.7E5 1100 570
TNN L 50000 0 411 550 1100 1800 3.4E5 1600 630
TNN S 500 10 412 770 340 1200 66000 240 630
TNN S 50000 10 413 640 1100 2300 2.2E5 1200 720
TNN L 50000 10 414 610 1100 2300 2.2E5 1200 800
TNN L 500 0 415 8900 1100 6300 3.2E5 810 930
ELM L 50000 10 416 9200 1000 54000 2.6E9 520 44000
ELM L 50000 0 417 9600 1100 56000 2.6E9 550 49000
ELM S 50000 10 418 9200 1000 54000 1.8E8 520 1.9E6
ELM S 50000 0 419 9700 1100 57000 1.9E8 550 2.0E6